RC meeting and Run Plan, Friday Sept. 22, 2023

From SBS wiki
Jump to: navigation, search

Return to the RC Meeting Calendar

RC Meeting Information

Time: Monday - Friday at 3:30pm

  • Location: Counting house conference room
  • ZoomGov Meeting Connection Information


Useful Links


Hall A call-in instructions:

  • Call RC first.
  • Techs on site 7-4:30 Mon-Th, 7-11AM Fr. RC will call Hall A work coordinator (Zach) during normal working hours. After hours, RC calls Tech on-call.
  • After 9pm everything non-emergency needs to be deferred. What is emergency and what's not? Call RC. RC talks with Mark to decide.

Hall Access procedure:

  • Access check list (binder on the counting room desk) needs to be filled out before making an access.
  • Scan (take pictures with phone) and post it to the logbook.
  • Any work inside the SBS, BB boundary need to have PS locked out -- requires two QEWs to throw the big switch. Exception: field measurements.

New instruction for shift crew when changing target:

  • Beam restoration instructions have been updated. Printed version is in the counting room.
  • When we change target, make sure the SOFs of Target BPMs are set to the correct values.
  • SOFs for different targets: Needs an update after restoring beam tonight (Note: SOFs for "No Target" has been updated and they are same as the He3 target values)

Status Update

Overall Summary

  • Thurday Swing (09/21/2023): Moeller measurement was completed around 9pm after the DC power fixed the issue related to Moller dipole magnet. During the time DC power were fixing dipole (took ~ 4hr and completed around 5pm), accelerator was also down for the maintenance of 1L11 transformer wiring https://logbooks.jlab.org/entry/4184663 . So beam was available to the hall only after 7 pm.
Polarisation for HWP OUT: 80.61+/-0.24 and HWP IN: -80.39+/-0.21 https://logbooks.jlab.org/entry/4184594

After the Moeller, we followed the beam checkout procedure and took a carbon hole run: https://logbooks.jlab.org/entry/4184795 which looks beam centered to the hole. So, reference cell was purged and filled with H2: https://logbooks.jlab.org/entry/4184808 and started taking data on H2 reference cell with correct SOFs.

  • Friday OWL (09/22/2023): Beam current was ramped up to 30 uA, continued taking H2 data on reference cell. We took around 4.5M events on H2, and then switch to He3 target towards the end of owl shift.

Accelerator status

Events collected :

  • H2 data: 4.5M events (OWL shift)
  • He3 data: 0.26 C (corrected charge)

Hall A Subsystems Status


  • Polarization: 32.6% (2:20 pm)



Requires opportunistic hall access

  • Cooling jet hardware limit (Todd/Ellen)
  • SBS GEMs made DAQ unstable (GEM group)
  • HCAL/BBCALThreshold GUI readback (Alex/Jiwan/Ellen)

Shift schedule

  • Urgent: Need a Shift leader for tomorrow Swing shift and Tuesday Day shift


Run plan for the weekend

  • Production run on polarized He3 target @35uA beam current (?? can we go to 40uA if polarization is greater than 35%??)

List of items in case target isn’t ready:

  1. Take data on reference cell at 10 atm H2 (portions that add up to 1 day)
  2. BigBite Sieve slit => only during day (Techs needed)
    1. with 8 carbon foils and take optics for 1 hr at 40 uA (Sean/Andrew - confirm time)
    2. 1 hour for single foil (magnets off) => Chance BBcal HV setting to appropriate values
  3. BigBite/single foil location reconstruction (~20 mins?) - BigBite + HCAL detector experts on call consulting analysis experts like Andrew/Sean (magnet on)